OpenStudy (anonymous):

Jessica read 3/5 of a book in 3 days. On the first day she read 1/4 of the book and on the 2nd day she read 1/6 of the book. What fraction of the book did she read on the 3rd day?

OpenStudy (callisto):

3/5 - 1/4-1/6 = 11/60

OpenStudy (anonymous):

What fraction of the book did she read on the 3rd day, = 3/5 - 1/4-1/6 = 11/60

OpenStudy (callisto):

you actually don't have to rewrite my answer lol

OpenStudy (anonymous):

haha, i know, but i solve it my self,. LOL

OpenStudy (anonymous):

callisto, help me with my problem.

OpenStudy (callisto):

that's good but you're not correct in copying 'What fraction of the book did she read on the 3rd day' you are answering not asking so you need to make a change

OpenStudy (anonymous):

oh yeah, i noticed that, maybe because im tired of typing, LOL